### v1

@@ Abstract
Mixture of Experts (MoE) LLMs have recently gained attention for their ability to enhance performance by selectively engaging specialized subnetworks or “experts” for each input. However, deploying MoEs on memory-constrained devices remains challenging, particularly when generating tokens sequentially with a batch size of one, as opposed to typical high-throughput settings involving long sequences or large batches. In this work, we optimize MoE on memory-constrained devices where only a subset of expert weights fit in DRAM. We introduce a novel cache-aware routing strategy that leverages expert reuse during token generation to improve cache locality. We evaluate our approach on language modeling, MMLU, and GSM8K benchmarks and present on-device results demonstrating 2$\times$ speedups on mobile devices, offering a flexible, training-free solution to extend MoE’s applicability across real-world applications.
[Uncaptioned image][Uncaptioned image]
Figure 1: Overview of our proposed cache-aware routing method. (Left) Our proposed method enhances the cache hit-rate of existing MoEs while retaining the original model accuracy by incorporating a cache prior. (Right) Throughput of our caching method for the 4-bit and 8-bit quantized Qwen1.5-MoE models deployed on two mobile devices with 10GB and 16GB available memory and cache size of 30 and 45 experts, respectively. Our proposed Cache-Aware Routing method significantly enhances the token generation throughput compared to the baseline leveraging a Least Recently Used (LRU) Cache. https://en.wikipedia.org/wiki/Universal_Flash_Storage
footnotetext: 1Anonymous Institution, Anonymous City, Anonymous Region, Anonymous Country. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
@@ 1 Introduction
Mixture of Experts (MoE) models have emerged as a powerful approach in large language models (LLMs), offering advantages in scalability and efficiency Fedus et al. (2022b; a). By selectively activating a subset of experts for each input, MoEs handle diverse data distributions and capture complex patterns more effectively than traditional dense models. Recent models like GPT-4 (OpenAI et al., 2024), Gemini (Team et al., 2024), and Mixtral (Jiang et al., 2024) provide evidence for MoEs’ success in leveraging specialized sub-networks to achieve superior performance.
Gemini (Team et al., 2024), and Mixtral (Jiang et al., 2024) provide evidence for MoEs’ success in leveraging specialized sub-networks to achieve superior performance. Simultaneously, small language models (SLMs) with MoE architectures, such as OLMoE Muennighoff et al. (2024), Phi-3.5-MoE Abdin et al. (2024), and Qwen-MoE Qwen Team (2024), have shown promise for deployment on memory-constrained devices due to their ability to maintain high performance with fewer active parameters per token.
@@
The sheer size of these models often surpasses the available DRAM capacity, and loading model weights directly from flash memory is detrimental to latency. To mitigate this, DRAM can be used as a cache to store a subset of expert weights, but this approach is only effective if the cache hit rate is high. Ideally, MoE models should exhibit high cache locality, where a small subset of experts is repeatedly accessed over extended sequences of tokens. Addressing these challenges requires innovative caching-friendly expert routing strategies to manage memory effectively during inference while maintaining the performance benefits of MoE architectures.
@@
Previous works have attempted to enhance MoE model efficiency by pre-fetching expert layers using hidden states from prior layers and employing Least Recently Used (LRU) caching to store recently used experts Eliseev & Mazur (2023); Yuan et al. (2024). However, these methods rely on assumptions about expert locality and predictability, which can be unreliable. A major issue is the lack of temporal locality in the expert selection of state-of-the-art (SoTA) MoEs, leading to inefficient LRU caching due to frequent evictions and reloads, resulting in a poor cache hit rate.
@@
In this work, we aim to utilize already-trained, high-performing MoE models by integrating cache locality priors to enhance their suitability for memory-constrained devices. Specifically, we propose methods to improve expert reuse during token generation, thereby increasing the expert cache hit rate. This approach is training-free and can be applied directly to off-the-shelf MoE models, enhancing their efficiency for on-device deployment.
@@
To enable efficient inference on memory-constrained devices, we first analyze the temporal locality in expert selection across four SoTA MoEs, identifying factors that contribute to poor cache hit rates. We demonstrate their sensitivity to expert dropping and random expert swapping in Section 2.3. To address these issues, we propose manipulating router logits to increase the probability of selecting cached experts, eliminating the need for additional training. In summary, this work makes the following contributions:
  • We demonstrate that SoTA MoE models are not optimized for cache efficiency, exhibiting poor latency performance when an LRU cache is implemented. Additionally, we find that these models are less affected by permutations in the selection of lower-ranked experts, indicating the feasibility of lossy routing strategies.
@@
  • We propose a method to significantly improve the cache hit rate of experts, enhancing the throughput for token generation in existing SoTA MoEs without any training. Our approach enables balancing the trade-off between cache hit rate (and thus token generation latency) and model accuracy under a fixed memory budget.
@@
  • We report the performance of the proposed routing strategy for language modeling and present evaluation results on MMLU and GSM8K benchmarks. Surprisingly, we show that imposing some degree of routing consistency can enhance benchmark scores while making these models more on-device friendly.
@@
  • We present on-device results of our proposed approach on mobile devices with various memory constraints, demonstrating a consistent speed-up of up to 2×2$\times$2 × compared to the regular routing with LRU caching. Due to being training-free, our approach can be deployed across a range of real-world scenarios and diverse hardware configurations
@@ 2 Background and Motivation
In this section, we provide an overview of Mixture of Experts (MoE) models, followed by a sensitivity analysis that highlights the impact of expert selection on model performance.
@@ 2.1 Preliminaries of MoE
The sparse MoE layer consists of a set of $N$ expert networks $\{E_{1},E_{2},\ldots,E_{N}\}$, and a routing network $G$ which selects a subset of these experts based on their relevance to the input tokens. Let $\mathbf{x}\in\mathbb{R}^{d}$ denote the input token. The output of each MoE layer during inference is given by:
@@
$\mathbf{w}$ $=\sigma(\mathbf{z})=\sigma(G(\mathbf{x}))$ (3)
where $\mathbf{z}=G(\mathbf{x})$ denotes the logits assigned by the router to the experts, $\sigma$ denotes the softmax operation, and $\mathbf{r}$ is an expert ranking vector based on the weights, of which only the Top-K experts are selected. Note that the weights assigned to each expert might be re-normalized in some cases after the Top-K selection in Equation 1.
In conventional MoEs, the Top-K routing mechanism to select a subset of experts is inherently dynamic and input-dependent, but can lead to significant variability in routing decisions from one token to the next. This lack of consistency poses a fundamental challenge for achieving high cache hit-rates, as it reduces the likelihood that the selected experts for the given token are already present in cache. This can be seen in Table 2, where cache lifetimes are very short, and cache miss rates are high for all models deploying a simple LRU cache policy.
@@ 2.2 MoE deployment on memory-constrained devices
Deploying MoEs on memory-constrained devices, such as mobile phones, presents unique challenges due to limited memory resources. As shown in Figure 1 (Left), these devices typically feature a combination of DRAM and flash storage, each with distinct characteristics regarding speed and capacity. DRAM provides high bandwidth but limited capacity, while flash storage offers larger capacity at the expense of slower access speeds.
Figure 2: Expert sensitivity analysis. We show the effect of dropping or replacing experts as selected by the router. The x-axis represents the expert rank (ordered by their scores) and the y-axis shows the Wikitext validation perplexity (lower values indicate better performance). The left figure illustrates the effect of dropping all experts ranked higher than $k$, whereas the right figure depicts the impact of randomly replacing the expert at rank $k$.
In the context of MoE models, not all parameters can fit into the limited DRAM available. Static parameters, such as attention weights that do not change during inference, can be stored permanently in DRAM. In contrast, given the dynamic nature of expert selection, only a subset of experts is loaded into DRAM at any given time. This subset acts as a cache, enabling faster access to frequently/previously used experts.
A high cache hit rate is important because it means experts are quickly available from DRAM, reducing the overall latency. The cache hit rate measures the proportion of selected experts that are already present in the cache at the time of selection. Let the cache set $C$ denote the indices of the experts that are currently loaded into the DRAM cache, recall that $\mathbf{r}[:K]$ are the selected experts by the routing network. The cache hit rate is then defined as:
${\operatorname{hit-rate}}=\frac{{\operatorname{count}}(\mathbf{r}[:K]\cap C)}{K}.$ (4)
The $\operatorname{miss-rate}$ is simply $1-\operatorname{hit-rate}$.
A major motivation of our work is to enhance the cache hit rate of experts without compromising the original predictive performance of the model.
To achieve this, we design various functions to update the ranking vector $\mathbf{r}$ used to select the Top-K experts without making any other changes to the MoE forward pass as shown in Equation 1. This allows for a model-agnostic method that can be universally applied.
Our goal is to find new ranking vectors promoting experts that are currently in cache, while minimizing the effect on downstream model performance.
@@ 2.3 MoEs are less sensitive to dropping experts with lower scores
The potential for MoE models to become more cache-friendly without sacrificing task performance hinges on their robustness to variations in expert selection. To evaluate this assumption, we conduct a series of analyses on four different MoE architectures: DeepSeek-V2-Lite DeepSeek-AI et al. (2024), Qwen1.5-MoE-A2.7B Qwen Team (2024), Phi-3.5-MoE Abdin et al. (2024), and Mixtral-8x7B Jiang et al. (2024). Our findings demonstrate that these models can tolerate some deviations from the router-selected experts as long as the experts with the highest weights are kept.
The left plot in Figure 2 illustrates the impact of completely removing (pruning) experts ranked at or above a specified index. Notably, the removal of the second highest ranked expert leads to a significant performance drop across all four models. However, for models with a larger number of active experts, such as Qwen and DeepSeek, performance quickly recovers, allowing for the pruning of higher-ranked experts with minimal performance loss.
To illustrate the significance of expert rank while maintaining the total number of activated experts, we replaced the expert ranked $k$ with a randomly selected expert, as shown in the right plot of Figure 2. The results indicate that substituting the most important expert (rank one), which has the highest score, drastically affects model performance. Replacing the second ranked expert also leads to severe degradation, but for the higher ranks, the effect of swapping experts is minimal.
This sensitivity analysis of MoEs shows that dropping experts from the Top-K selection can adversely affect model accuracy and efficiency. In particular, all models are sensitive to dropping of the Top-1 and Top-2 experts. However, the results also suggest that while retaining the highest-ranked experts is crucial, most MoE models are relatively robust to minor variations in the experts chosen by the router.
@@ 3 Cache-aware expert routing
In this section, we present our primary techniques designed to enhance the consistency of routing decisions during token generation, which in turn improves the expert cache hit rate and overall model throughput. We propose a general training-free method that can be applied to off-the-shelf MoE models to improve their efficiency for on-device deployment. Additionally, we introduce two simple baselines to increase cache hit rates: Max Rank and Cumulative Probability Threshold.
Figure 3: Our proposed cache-aware routing method. A bias is added to the logits to promote the selection of in-cache experts.
@@ 3.1 Max Rank Approach
In Section 2.3 we demonstrated that MoEs exhibit some robustness to the swapping of their experts with random ones. Building on this insight, we aim to deviate from the Top-K set of experts $\mathbf{r}[:K]$, selected by the router and instead promote experts that are already present in the cache $C$. We first define a general promotion operation that elevates the ranking of all experts in an ordered subset $\mathbf{r}^{\text{subset}}$:
@@
Here $\oplus$ denotes the concatenation operation and the set subtraction $\setminus$ preserves the order of it’s left operand.
@@
Finally, the pseudocode for the max-rank algorithm with always keeping the Top-J experts is shown in Algorithm 1.
@@ 3.2 Cumulative Probability Threshold Approach
We address this issue in the cumulative probability threshold approach by dynamically choosing the max-rank $M$ for every layer and every input $\mathbf{x}$. We determine $M$ by summing the sorted probabilities of the router outputs $G(\mathbf{x})$ from highest to lowest until a predefined cumulative probability threshold $p$ is reached:
$M=\operatorname{min}i\ \ s.t.\sum_{j=1}^{i}G(\mathbf{x})[r_{i}]\geq p.$ (8)
Using the set of $M$ values found for each layer and each input, we apply the max-rank strategy from Equation 7 to select a new set of experts. The pseudocode for the cumulative probability threshold approach can be found in Algorithm 2. This method is more dynamic as it accounts for the uncertainty in the router’s predictions. When the set of experts required to reach the threshold is larger, it indicates less confidence from the router, allowing for more flexibility in replacing a non-cached expert with the highest probability expert available within this set that is already in the cache.
@@
3.3 Cache Prior Reranking

### v2

@@ Abstract
Mixture of Experts (MoE) LLMs enhance performance by selectively activating specialized subnetworks (“experts”) per input. While MoEs offer efficiency benefits through distributed inference in typical high-throughput settings, deploying them on memory-constrained devices remains challenging, particularly for sequential token generation with batch size one. In this work, we optimize MoE for such constrained environments, where only a subset of expert weights fit into DRAM. Through empirical analysis, we show MoEs can tolerate careful deviations in expert selection with minimal predictive performance loss. Inspired by this observation, we propose a novel cache-aware routing strategy that leverages expert reuse during token generation to significantly improve cache locality. Evaluating on language modeling, MMLU, and GSM8K benchmarks, our method reduces cache miss rates by over 50%, with negligible impact on perplexity (0.1%–3%) and downstream task accuracy (<0.1%). Unlike prior methods limited by the optimal oracle cache bound, our approach surpasses this theoretical limit by allowing slight flexibility in expert selection. Finally, we present on-device results demonstrating 2$\times$ speedups on mobile hardware, offering a flexible and training-free solution to extend MoE’s applicability across real-world applications.
@@ 1 Introduction
Mixture of Experts (MoE) models have emerged as a powerful approach in large language models (LLMs), offering advantages in scalability and efficiency (Fedus et al., 2022b; a). By selectively activating a subset of experts for each input, MoEs handle diverse data distributions and capture complex patterns more effectively than traditional dense models. Recent models such as DeepSeek-V3 (DeepSeek-AI et al., 2024), Qwen2.5-Max (Yang et al., 2024), GPT-4 (OpenAI et al., 2024), Gemini (Team et al., 2024), and Mixtral (Jiang et al., 2024) provide evidence for MoEs’ success in leveraging specialized sub-networks to achieve superior performance.
Gemini (Team et al., 2024), and Mixtral (Jiang et al., 2024) provide evidence for MoEs’ success in leveraging specialized sub-networks to achieve superior performance. Simultaneously, small language models (SLMs) with MoE architectures, such as OLMoE (Muennighoff et al., 2024), Phi-3.5-MoE (Abdin et al., 2024), and Qwen-MoE (Qwen Team, 2024), have shown promise for server deployment due to their ability to maintain high performance with fewer active parameters per token.
Figure 1: Overview of our proposed cache-aware routing method. (Left) MoE models are hosted in slower Flash storage due to their size, with only a subset of expert weights cached in faster DRAM. Our cache prior method is cache-aware and adjusts expert selection to promote experts already in DRAM, significantly reducing cache misses and improving inference efficiency. (Right) Throughput of our routing method for the 4-bit and 8-bit quantized Qwen1.5-MoE models deployed on two mobile devices with 12GB and 16GB available memory and cache size of 30 and 45 experts, respectively. Our proposed Cache-Aware Routing method significantly enhances the token generation throughput compared to the baseline leveraging a Least Recently Used (LRU) Cache. https://en.wikipedia.org/wiki/Universal_Flash_Storage
@@
These models often exceed available DRAM, and loading model weights directly from flash storage is detrimental to latency. To mitigate this, DRAM can be used as a cache for expert weights, but this approach is only effective if the hit rate is high. Ideally, MoE models should have strong cache locality, with a small set of experts being accessed repeatedly over long token sequences. This requires better strategies for expert selection and cache optimization during inference.
@@
Previous work has tried to improve MoE efficiency by pre-fetching experts based on hidden states from earlier layers and using LRU caching to store recently used experts (Eliseev & Mazur, 2023; Yuan et al., 2024a). However, these methods assume that expert selection follows predictable patterns and has a strong locality. Temporal locality, in the context of MoE caching, refers to the principle that recently accessed experts are likely to be accessed again in subsequent token generations. A key challenge, which we analyze empirically in Section 4.6, is that state-of-the-art MoEs often lack strong temporal locality in expert selection, leading to inefficient Least Recently Used (LRU) caching strategies due to frequent evictions and reloads, which results in a poor cache hit rate.
@@
In this work, we aim to utilize already-trained, high-performing MoE models by incorporating cache locality priors to make them more efficient for memory-limited devices. Our method increases expert reuse during token generation, improving the cache hit rate. This approach is training-free and can be applied directly to off-the-shelf MoE models, enhancing their efficiency for on-device deployment.
@@
To enable efficient inference on memory-constrained devices, we first analyze the temporal locality in expert selection across four SoTA MoEs. We demonstrate their sensitivity to expert dropping and random expert swapping in Section 2.3, which shows the potential for exploring alternative routing strategies. Motivated by these observations, we propose a method that manipulates router logits to increase the probability of selecting cached experts, all without the need for additional training. In summary, this work makes the following contributions:
  • We demonstrate that state-of-the-art MoE models lack temporal consistency in expert selection, which leads to suboptimal latency performance when caching strategies are employed. However, our findings indicate that these models exhibit low sensitivity to variations in the selection of lower-weighted experts, suggesting that lossy routing strategies can be implemented without substantial degradation in overall performance.
@@
  • We propose a method to significantly improve the cache hit rate of experts, enhancing throughput for token generation in existing MoEs without requiring additional training. Our approach strikes a balance between cache hit rate (and token generation latency) and model accuracy under a fixed memory budget.
@@
  • We report the performance of the proposed routing strategy for language modeling and present evaluation results on MMLU and GSM8K benchmarks. Surprisingly, we show that imposing some degree of routing consistency enhances benchmark scores while making these models more on-device friendly.
@@
  • We present on-device results of our proposed approach on two different mobile devices with various memory constraints, demonstrating a consistent speed-up of up to 2×2$\times$2 × compared to the original routing with LRU caching. Our training-free method is adaptable and can be deployed across a variety of real-world scenarios and diverse hardware configurations, making it suitable for devices with different memory limitations.
@@ 2 Background and Motivation
In this section, we give an overview of MoE models, followed by a sensitivity analysis that looks at how expert selection affects model performance, showing the potential for alternative routing strategies explored in the next section.
@@ 2.1 Preliminaries of MoE
The sparse MoE layer consists of $N$ expert networks ${E_{1},E_{2},\ldots,E_{N}}$ and a routing network $G$ that selects a subset of experts based on their relevance to the input token $\mathbf{x}\in\mathbb{R}^{d}$. The output of the MoE layer during inference is given by:
@@
$\mathbf{w}=\sigma(\mathbf{z})=\sigma(G(\mathbf{x})).$ (3)
Here, $\mathbf{z}=G(\mathbf{x})$ represents the logits assigned by the router to the experts, $\sigma$ is the softmax function, and $\mathbf{r}$ is a ranking vector of experts based on the weights, from which only the top-$K$ experts are selected. The weights assigned to each expert may be re-normalized after the top-$K$ selection in Equation 1.
In conventional MoEs, the top-$K$ routing mechanism for selecting experts is dynamic and input-dependent, which can cause significant variability in the experts chosen for each token. This inconsistency makes it difficult to achieve high cache hit rates, as it reduces the chance that selected experts are already in the cache. As shown in Table 10, cache lifetimes are short, and cache miss rates are high for all models using a simple LRU caching policy.
@@ 2.2 MoE Deployment on Memory-Constrained Devices
Deploying MoEs on memory-constrained devices, like mobile phones, is challenging due to limited memory resources. As shown in Figure 1 (Left), these devices typically feature a combination of DRAM and flash storage, each with different speed and capacity characteristics. DRAM offers high bandwidth but limited capacity, while flash storage provides larger capacity at the cost of slower access speeds.
Figure 2: Expert sensitivity analysis. We show the effect of dropping or replacing experts as selected by the router. The x-axis represents the expert rank (ordered by their scores) and the y-axis shows the Wikitext validation perplexity (lower values indicate better performance). The dashed lines represent the baseline perplexity of the MoE models.
The left figure illustrates the effect of dropping all experts ranked higher than $k$, whereas the right figure depicts the impact of randomly replacing the expert at rank $k$.
MoE models are typically larger than dense models, therefore, not all parameters may fit into the limited DRAM available on many devices. Static parameters, such as attention weights that do not change during inference, can be stored permanently in DRAM. In contrast, due to the dynamic nature of expert selection, only a subset of experts is loaded into DRAM at a time. To improve efficiency, recently used experts can be retained in a cache, reducing the need to reload them from slower storage. For an effective caching strategy, a high cache hit rate is important because it means experts are quickly available from DRAM, reducing the overall latency. The cache hit rate is defined as the proportion of selected experts that are already in the cache at the time of their selection. Let the cache set $C$ denote the indices of experts currently stored in DRAM, and $\mathbf{r}[:K]$ be the experts selected by the routing network. The cache hit rate can be expressed as:
${\emph{hit\text@underline{\phantom{ }}rate}}=\frac{{\operatorname{count}}(\mathbf{r}[:K]\cap C)}{K}.$ (4)
[…]
The goal of this work is to improve the expert cache hit rate without affecting the model’s predictive performance. To achieve this, we propose functions that modify the ranking vector $\mathbf{r}$ used to select the top-$K$ experts without making any other changes to the MoE forward pass as shown in Equation 1. Our objective is to create new ranking vectors that prioritize experts already in cache while minimizing any negative impact on model performance. This allows for a model-agnostic method that can be universally applied.
@@ 2.3 MoEs Are Less Sensitive to Dropping Experts with Lower Scores
A key opportunity for making MoE models more cache-friendly without compromising task performance is understanding how sensitive the models are to changes in expert selection. To explore this, we analyze four different MoE architectures: DeepSeek-V2-Lite (DeepSeek-AI et al., 2024), Qwen1.5-MoE-A2.7B (Qwen Team, 2024), Phi-3.5-MoE (Abdin et al., 2024), and Mixtral-8x7B (Jiang et al., 2024). Our results show that these models can tolerate deviations in the expert selection, as long as the experts with the highest weights are retained.
The left plot in Figure 2 illustrates the impact of completely removing (pruning) experts ranked at or above a specified index on Wikitext perplexity. Removing the second highest ranked expert leads to a performance drop across all four models. However, models with more active experts, like Qwen and DeepSeek, recover quickly, allowing higher-ranked experts to be pruned with minimal performance loss.
To examine how the rank of an expert impacts performance while keeping the total number of active experts constant, we replaced the expert ranked $k$ with a randomly selected expert, as shown in the right plot of Figure 2. The results show that swapping the top-ranked expert severely compromises model performance. Replacing the second-ranked expert also causes noticeable degradation, but the model remains functional. Beyond this, in granular MoEs (with a larger number of smaller active experts), the model becomes highly resilient to expert selection changes. This suggests that while top-$1$ is critical for all MoEs, the top-$2$ expert in standard MoEs such as Mixtral and Phi-MoE can be swapped without a major loss in perplexity. Granular models maintain flexibility in expert selection and incur minimal performance loss by swapping the 3rd highest ranked expert and beyond.
Interestingly, further analysis (Appendix A) using a greedy search to find the optimal expert combination suggests that the router’s predictions are often suboptimal. For Mixtral-8x7B, the router’s predicted top-$2$ experts yield the best performance only 28% of the time on average with a maximum of 38% in the last layer. This supports the idea that lower-ranked experts can be replaced with little impact on performance. Taken together, these findings highlight that while preserving the highest-ranked experts is crucial, MoE models offer flexibility in expert selection—an insight we leverage throughout this work.
@@ 3 Cache-aware Expert Routing
In this section, we introduce our main method along with two simple baselines: Max Rank and Cumulative Probability Threshold. These methods, progressively more refined, balance cache hit rate with downstream task performance, improving model throughput and efficiency. All of our methods are general and training-free, making them easily applicable to off-the-shelf MoE models for on-device deployment.
@@ 3.1 Max Rank Approach
In Section 2.3, we demonstrated that MoEs exhibit some robustness to the swapping of their experts with random ones. Building on this insight, we aim to deviate from the standard top-$K$ set of experts $\mathbf{r}[:K]$, selected by the router and instead promote experts that are already present in the cache $C$. We first define a general promotion operation that elevates the ranking of all experts in an ordered subset $\mathbf{r}^{\text{subset}}$:
@@
Here $\oplus$ denotes the concatenation operation and the set subtraction $\setminus$ preserves the order of it’s left operand. Importantly, all sets in this context are ordered sets and maintain the order of elements as in their original ranking. This ensures that the promotion operation is well-defined, as it preserves the relative ordering of experts throughout.
@@
Finally, the pseudocode for the max-rank algorithm with always keeping the top-$J$ experts is shown in Algorithm 1. Additionally, we provide an intuitive explanation of the algorithm with an example in Appendix B.
@@ 3.2 Cumulative Probability Threshold Approach
We address this issue in the cumulative probability threshold approach by dynamically choosing the max-rank $M$ for every layer and every input $\mathbf{x}$. The process is conceptually similar to the sampling approach by Holtzman et al. (2020) in which tokens are sampled from a dynamic set containing the vast majority of the probability mass. We determine $M$ by summing the sorted probabilities of the router outputs $G(\mathbf{x})$ from highest to lowest until a predefined cumulative probability threshold $p$ is reached:
$M=\operatorname{min}i\ \ s.t.\sum_{j=1}^{i}G(\mathbf{x})[r_{j}]\geq p$ (8)
Using the set of $M$ values found for each layer and each input, we apply the max-rank strategy from Equation 7 to select a new set of experts. We provide an schematic explanation of this approach in Appendix B. The pseudocode for the cumulative probability threshold approach can be found in Algorithm 2. This method is more dynamic as it accounts for the confidence in the router’s predictions. When the set of experts required to reach the threshold is larger, it indicates less confidence from the router (flat distribution), allowing for more flexibility in replacing a non-cached expert with the highest probability expert available within this set that is already in the cache.
@@
3.3 Cache-Prior Reranking